what is the answer for these problems. Iam having diffcult with just these, please show work and use sub. or eilm. thanks.
1. x-9y=237
5x+3y=-267

2. x-4y=119
16y--432

3. 4x+5y=1
14x+5y=191

4. 8x+y=74
3x+19y=363

5.9x+17y=-475
7x+11y=-285

6. 9x+13y=536
9x+16y=587

7. 10x-y+-145
13x+3y=-339

8. 17x-8y+=154
17x-12y+=150

9. 11x-4y=364
2x-19y=322

10. x+2y=47
x-3y=7

11. 3x-4y=-198
x-3y=-111


12. 12x+13y=491
12x+19y=629



Answer by elima(1433)

I will solve two problems, one with substitution and one with elimination.
1)x-9y=237
5x+3y=-267
Substitution;
First of all solve one of the eq. for one of the variables; We will solve the first one for x;
x-9y=237
x=237+9y
Now substitute that for x in the other eq.
5(237+9y)+3y=-267
1185+45y+3y=-267
48y=-267-1185
48y=-1452
y=-30 1/4
Now that we have y plug that answer into one eq. and solve for x;
5x+3(-30 1/4)=-267
5x-90 3/4=-267
5x=176 1/4
x=35 1/4
So;
x=35.25
y=-30.25
Now we will solve this by elimination;
3)4x+5y=1
14x+5y=191
First of all we need to multiply eq.2 by -1;
(-1)14x+5y=191
-14x-5y=-191
Now we will add eq.1 and eq.2 together;
4x+5y=1
-14x-5y=-191
-------------
-10x=-190
x=19
Now we have x, so plug that into one of the eq. to get y;
4(19)+5y=1
76+5y=1
5y=-75
y=-15
So ;
x=19
y=-15
Try and work out the others by yourself. If you have trouble let me know.
=)